OpenStudy (anonymous):

What are the solutions of the equation x4 + 3x2 + 2 = 0? Use u substitution to solve.

OpenStudy (gorv):

let x^2=t t^2+3t+2=0 t^2+2t+t+2=0 t(t+2)+(t+2)=0 (t+2)(t+1)=0 t=-2,-1 x^2=-2 x=i*2^(1/2) x=i

OpenStudy (gorv):

@sexy12